# PIR: Photometric Inverse Rendering with Shading Cues Modeling and Surface Reflectance Regularization

### v2

@@ 3.1 Overview
Our pipeline commences with estimating the object’s geometry and surface diffuse albedo using the off-the-shelf neural surface reconstruction framework, NeuS [42]. Subsequently, we utilize physics-based rendering to jointly refine the geometry and materials of the object as well as the position and intensity of the flashlight. Our approach leverages differentiable rendering techniques to accurately model self-shadows and indirect illumination in photometric images, thereby achieving robust material decomposition. Additionally, we reduce the ambiguities of surface reflectance decomposition by integrating self-supervised DINO [9] features from multi-view images. Our method can export the mesh and texture maps of the optimized 3D models, which can be seamlessly integrated into conventional rendering pipelines.
@@ 3.3 Light Source Optimization
$\boldsymbol{w}_{i}(\boldsymbol{x})=\frac{(\boldsymbol{o}+\Delta_{o})-\boldsymbol{x}}{\|(\boldsymbol{o}+\Delta_{o})-\boldsymbol{x}\|^{2}},L_{i}(\boldsymbol{w}_{i};\boldsymbol{x})=\frac{L}{\|(\boldsymbol{o}+\Delta_{o})-\boldsymbol{x}\|^{2}},$ (4)
@@ 4.2 Comparison with Existing Methods
For a fair comparison of material decomposition, we adapted the physical shader in WildLight to roughplastic model [61], as utilized by Mitsuba. We then integrated the decomposed material components along the rays to generate the material maps of the view. Comparison with the methods on volume rendering method DRV [5] and the mesh-based approach PSDR [40] were not conducted in our study, as their codes are not available. We also compared with other implicit methods presented by [82], which recover neural fields while considering inter-reflections.
@@
Figure 6: Quantitative and Cisual Results of Light Optimization.
@@ 4.4 Ablation Studies
The synthetic data rendered for the ablation study on inter-reflection is in a collocated camera-lighting setting to overview self-shadows and different physical shader settings used in different methods. We ablate the inter-reflection calculation and compare the results in Table 3 and Fig. 7. Our method accurately estimates the materials, and without the inter-reflection modeling, the predicted specular albedo has artifacts, and diffuse albedo baked the indirect illumination.
Figure 9: Relighting Results. Our method supports PBR relighting with the estimated materials.
@@ 5 Conclusion
In this paper, we present a new and effective inverse rendering approach for reconstructing object shapes, materials, and lighting from photometric images. Our method optimizes the light source position to account for self-shadows and employs an online strategy for modeling inter-reflections through a differentiable rendering layer. Additionally, we incorporate the DINO regularization to help the decomposition of surface reflectance. Extensive experiments on synthetic and real datasets demonstrate that our method can address misalignments between camera and light sources and surpass state-of-the-art methods in material decomposition. Thanks to its ability to precisely compose objects, DINO regularization holds promise for broader applications in inverse rendering, including scenarios involving environmental illumination.
@@
Our method has several constraints. Primarily, we overlook the consistency of novel views captured by a moving flashlight in the geometry initialization stage. Secondly, the BRDF model we employ is tailored for solid reflective surfaces, thus our approach is not suitable for reflective surfaces. We leave these for our future work.
Acknowledgement This work was supported in part by NSFC with Grant No. 62293482, the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone. It was also partially supported by NSFC with Grant No. 62202409, the Shenzhen Science and Technology Program with Grant No. RCBS20221008093241052, the Shenzhen Outstanding Talents Training Fund 202002, the Guangdong Research Projects No. 2017ZT07X152 and No. 2019CX01X104, the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), and the Shenzhen Key Laboratory of Big Data and Artificial Intelligence (Grant No. ZDSYS201707251409055).

### v3

@@ 3.1 Overview
Our pipeline commences with estimating the object’s geometry and surface diffuse albedo using the off-the-shelf neural surface reconstruction framework, NeuS [43]. Subsequently, we utilize physics-based rendering to jointly refine the geometry and materials of the object as well as the position and intensity of the flashlight. Our approach leverages differentiable rendering techniques to accurately model self-shadows and indirect illumination in photometric images, thereby achieving robust material decomposition. Additionally, we reduce the ambiguities of surface reflectance decomposition by integrating self-supervised DINO [9] features from multi-view images. Our method can export the mesh and texture maps of the optimized 3D models, which can be seamlessly integrated into conventional rendering pipelines, as shown in Fig. 1.
@@ 3.3 Light Source Optimization
$\boldsymbol{w}_{i}(\boldsymbol{x})=\frac{(\boldsymbol{o}+\Delta_{o})-\boldsymbol{x}}{\|(\boldsymbol{o}+\Delta_{o})-\boldsymbol{x}\|},L_{i}(\boldsymbol{w}_{i};\boldsymbol{x})=\frac{L}{\|(\boldsymbol{o}+\Delta_{o})-\boldsymbol{x}\|^{2}},$ (4)
@@ 4.2 Comparison with Existing Methods
For a fair comparison of material decomposition, we adapt the physical shader in WildLight to roughplastic model [62], as utilized by Mitsuba. We then integrate the decomposed material components along the rays to generate the material maps of the view. Comparison with the methods on volume rendering method DRV [5] and the mesh-based approach PSDR [41] were not conducted in our study, as their codes are not available. We also compared with other implicit methods presented by [83], which recover neural fields while considering inter-reflections.
@@ 4.3 Results on Real Data
Figure 6: Quantitative and visual results of light optimization.
@@ 4.4 Ablation Studies
The synthetic data rendered for the ablation study on inter-reflection is in a collocated camera-lighting setting to avoid the influence of self-shadows and different physical shader settings used in different methods. We ablate the inter-reflection calculation and compare the results in Table 3 and Fig. 7. Our method accurately estimates the materials, and without the inter-reflection modeling, the predicted specular albedo has artifacts, and diffuse albedo baked the indirect illumination.
Figure 9: Relighting results with the estimated materials.
@@ 5 Conclusion
In this paper, we present an effective inverse rendering approach for reconstructing object shapes, materials, and lighting from photometric images. Our method optimizes the light source position to account for self-shadows and employs an online strategy for modeling inter-reflections through a differentiable rendering layer. Additionally, we incorporate the DINO regularization to help the decomposition of surface reflectance. Extensive experiments on synthetic and real datasets demonstrate that our method can address misalignments between camera and light sources and surpass state-of-the-art methods in material decomposition.
@@
Our method overlooks the consistency of novel views captured by a moving flashlight during geometry initialization, and the BRDF model is tailored for solid reflective surfaces. Future work will address these limitations and extend our approach to more complex imaging scenarios, including underwater environments [86, 25].
Acknowledgment
This work was supported in part by NSFC with Grant No. 62293482, the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone. It was also partially supported by NSFC with Grant No. 62202409, the Shenzhen Science and Technology Program with Grant No. RCBS20221008093241052, the Shenzhen Outstanding Talents Training Fund 202002, the Guangdong Research Projects No. 2017ZT07X152 and No. 2019CX01X104, the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), and the Shenzhen Key Laboratory of Big Data and Artificial Intelligence (Grant No. ZDSYS201707251409055).